Upstairs (non-receiver) Instrumentation Log

as of Sunday, 17-Dec-2017 17:38:46 AST

Go to Most Recent Entry


NameEmailDateComment
lswlwray@naic.eduTuesday, March 15, 2005 14:27:51Log Page created.
anollaanolla@naic.eduWednesday, March 16, 2005 16:07:13In order to switch between reading the rotary floor temperature and the rotary floor tilt sensor, added a relay to switch between the two. HP3488#1, bit 212. Voltmeter reads result.
ansanolla@naic.eduWednesday, May 25, 2005 13:18:34Now there are 3 recvers with its bias boxes tie to the UPS at the rf cabinet #2.These are the CB,CBH & XB. All of them are using indium phosphite rf amps at their Dewars , and in order to protect them from power dips they are tie to a UPS.
ansanolla@naic.eduTuesday, August 2, 2005 15:31:17The thru amps (UG-801D11) channel B input connector was misplaced . So Tim Hankings was not able to see this specific channel thru his system.
ansanolla@naic.eduWednesday, August 31, 2005 16:11:22The air compressor for the SBH & CBL shutters was not working. It was replace with a new one.
ansanolla@naic.eduThursday, September 1, 2005 11:37:41Tha HP meter 34970 who reads the cryo temp was replaced with the spare one.SN are US37037588 for the one installed & sn us3703689 for the bad one (given to LM to send it for repairs.)
ansanolla@naic.eduWednesday, September 14, 2005 13:50:14RF cable of pos 4 in the IF channel A sw broken , replaced.
ansanolla@naic.eduThursday, September 15, 2005 13:16:42The till sensor/Temperature meter relay swap network had an open ground in the till sensor mode.
ansanolla@naic.eduMonday, October 10, 2005 15:04:02This morning the rotary floor alarm was on all the time.So we call JR & LM to notify it. JR fix it up.
ansanolla@naic.eduThursday, October 13, 2005 10:36:27Yesterday we found that the Kepco 48 vdc power supply at the radar room was tripping off.We pinpoint that the emergency stop sw at the rotary floor had a short on it. Maybe the same problem that we had before that it was the 25DB feed thru rfi filter at the rotary floor Hoffman box.THESE SUPPLY IS VERY IMPORTANT TO THE OPERATION OF THE 430 MHZ RECEIVERS AT THE CH #1. It supplies the control voltage for several things , like the if selector sw for different receivers at the CH , LO swapping between systems like the CRR/LRR receivers L.O. , ect,ect.
ansanolla@naic.eduThursday, October 13, 2005 11:58:10Today JRB was checking the rotary floor emergency sw short and when he was checking the control pairs 53 & 54 the short ran away , try to duplicate it , but no way ; have to keep an eye on it.
ansanolla@naic.eduTuesday, October 18, 2005 12:49:51Dome rf cabinets doors were cleaned thru the rfi shielding stripes.The shielding stripes needs to be replaced in several parts specificlly in the rf cabinet 2 front door.
ansanolla@naic.eduMonday, October 24, 2005 10:50:33The Dome been weigthed so , we were not able to work in the Dome area due to this condition.
ansanolla@naic.comWednesday, October 26, 2005 13:54:41A protective cover was added to all the cryohead drivers in its power sw's to avoid accidental turn off.
ansanolla@naic.eduWednesday, February 1, 2006 11:25:57A third air conditioner unit added to the rotary floor,unit installed downhill.
ansanolla@naic.eduWednesday, February 15, 2006 13:11:48Air conditioner #3 frozeen , so call JV to check it.Meanwhile AC #1 turned on as replecement.
ansanolla@naic.eduWednesday, February 15, 2006 13:17:59Replacement He tank at the Dome just had .7 k instead 2.4 k. In the need of new He tank at the Dome.ANS will write a work order to lift it up.4 new tanks ordered thru Wanda.
ansanolla@naic.eduFriday, February 17, 2006 16:13:12The big vacuum pump at the Dome was left pumping in order to clean it out.It was contaminated by the SBH Dewar.
ansanolla@naic.eduMonday, February 20, 2006 11:30:55A full He tank (068594) raised to the Dome.
ansanolla@naic.eduThursday, March 9, 2006 12:36:26Jose Vives fix up the ALFA shutter , now when removing or instaling the shutter, the holding bracket will hold the shutter in place.
ansanolla@naic.eduThursday, March 23, 2006 16:22:05Remote AC sw installed back in cabinet 1 of rotary floor to control the VLBI calibrator. Bit 215 is the comand to control this unit. Also there is an spare AC sw whos control bit is 514.
ansanolla@naic.eduFriday, March 24, 2006 14:39:51The Dome remote AC sw power relocated at an UPS outlet , it wasn't.
ansanolla@naic.eduWednesday, April 5, 2006 11:40:16He tank at the Dome replaced.It was 717442 , now is 068594.
ansanolla@naic.eduMonday, April 10, 2006 12:51:06Shutters air compressor was not working.Check it and was found a loose connection. Anibal Rosado supplied an air line from their air compressor , so J Vives is working in the adapter of it to our system with quick lock air connectors.
ansanolla@naic.eduMonday, April 10, 2006 15:01:33Jose Vives installed the maintenance air compressor into our neumatic controlled shutters,our normal compressor is going to be used as back up.
ansanolla@naic.eduThursday, May 25, 2006 16:00:46Dome receivers cal driver unit had its power supply bad.
ansanolla@naic.eduMonday, June 12, 2006 14:45:25The former air compressor for the pneumatic shutters at the Dome is up again as an spare. It was fix by JVR.
ansanolla@naic.eduMonday, June 26, 2006 11:23:33Radar blanker antenna removed for maint.All rusted.
ansanolla@naic.eduFriday, July 7, 2006 11:29:56Ganesh told us to check the configuration of the Agilent sw driver 34970A on the rf cabinet at the Dome.The slot 1 is a 34908A 40 chann MUX , slot 2 is a 34907A MUX and slot 3 is free (Spare).
W. Genterwgenter@naic.eduFriday, July 7, 2006 16:10:48At 3:15 PM this afternoon, I was advised by Phil that the Power Meter was not communicating. Rey Velez and I went to the Gregorian. No key was found to open the tool box. I used my P-38 to open the shielded box in the right hand rack. The power meter was found in Stand-By. We turned on the meter and waited for the unit to display power levels. We replaced the cover. Phil called and indicated that he could not read the power meter again. We once again removed the shielded box cover and again the power meter was in Stand-By. We turned on the meter again. I noticed that the meter was very close to the closed position for the shielded box cover and pushed the power meter to the back about . Phil could still read the meter and was doing a TSYS. We attached the cover and the TSYS was continuing normally and did not stop as before. We closed the rack and returned to the control room. Everything is OK.
ansanolla@naic.eduThursday, July 13, 2006 10:50:12Cryo comps at the gimble on the Dome were without the secure bracket on four of them.So we installed the missing ones on them.
ansanolla@naic.eduThursday, July 13, 2006 11:05:51Cryo pump (small one) sn 445525 has a bad valve , it does not hold the vacuum. C. Rios is checking it.
ansanolla@naic.eduTuesday, August 1, 2006 11:31:06Spare air compessor for Dome's receiver shutters is ready to go in case that the maint people one fail.It has swage lock air connectors.
ansanolla@naic.eduThursday, August 10, 2006 11:02:57The remote cryo temp system was not working.So we found that rebooting the HP meter 39970 made the system comes alive again.ANS accidentally turn off the VI UPS for the 430 Mhz Mono driver , and walla for some reason that we don'tn know our shielded meter box was hooked into these UPS. Power for the shielded box swap to our Belkin UPS.
ansanolla@naic.eduMonday, October 9, 2006 14:50:16Dewar cryo temp monitor program was not running.Call downstairs in order to reset it.Now is ok.
ansanolla@naic.eduFriday, October 27, 2006 14:30:56The low frecuency converter was opened to check for an unstability on channel B at the 750 Mhz one. Ganesh found that there was an sma barrel on these converter with a lot of dirt on it,it was replaced with a new one , also all the connectors were cleaned & checked in both 750 mhz converters.
ansanolla@naic.eduMonday, October 30, 2006 13:12:58Channel B HP attenuator in the atennuator & power monitor chassis ( UG-081D16 )was bad , it had a 20 db loss thru it.
ansanolla@naic.eduThursday, November 2, 2006 11:47:45An rf cable was replaced in the channel B of the low frequency converter.It was the one who goes between the 300 Mhz converter (UG-801D09) out channel B and the IF sw 5B 2(UG-801D15).
ansanolla@naic.eduTuesday, November 28, 2006 09:15:08A new He regulator was installed to the helium tank at the Dome.Is a Victor Medalist P2500-500-580.It was installed on 11/15/2006.
ansanolla@naic.eduTuesday, February 20, 2007 15:06:12The rotary floor was cleaned out with the vacuum cleaner.
ansanolla@naic.eduMonday, April 30, 2007 13:39:32All systems were were wraped with plastic covers for protection during the painting project.RF cabinets were not.
ansanolla@naic.eduMonday, April 30, 2007 13:44:25The following chassis were taken down in order to replace toggle sw with lock up type.SBL bias chassis , Spare bias chassis , CBH bias & post amp chassis & the 430 MHz bias & post amp chassis.
Ganeshganesh@naic.eduMonday, February 4, 2008 14:09:292-12 GHz downconverter to 1.5GHz IF is not functional due to LO amp. failure. Users are advised to go thro' the 10 GHz upconverter path until replacement amp (from Miteq) is received. e.g. "ifsetup xb_10" instead of "ifsetup xb"
crvMonday, April 21, 2008 15:11:08Their was plobem with the shutters. The platform air compressor wasn't working so we installed the portable air compressor.
William Iguinawiguina@naic.eduTuesday, January 12, 2010 13:41:07Osvaldo y yo remplazamos la ventana del distomat #6, lectura anterior fue 8segundos a 1.5 de potencia, despues del cambio es de 3segundos a 3.5 de potencia.


Back to Top